find the vertex, line of symmetry , and maximum or minimum value of the quadratic function: f(x)= 2x^2 - 8x + 22
Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

The LOS is x = -b/2a
x = 8/4 = 2
x = 2 is the LOS
-----------------
The vertex is on the LOS
f(2) = 8 - 16 + 22 = 14
Vertex at (2,14)
---------
The vertex is the minimum

Answer by rfer(16322)

vertex (2,14)minimum
axis=2